OpenStudy (anonymous):

given -x+3y=11 find the parallel line that goes through the point (4,-2)

OpenStudy (anonymous):

y = (x/3) + (11/3) plug in (4,-2) -2 = (4/3) + b b = -3.33 y = (x/3) -3.33 try that

OpenStudy (nenadmatematika):

y=(1/3)x+11 is the form of the equation that you're given, your line has a same slope 1/3, so its equation is: y+2=1/3(x-4) so y=(1/3)x-10/3

OpenStudy (anonymous):

yea what he said

OpenStudy (anonymous):

thank you guys!

OpenStudy (nenadmatematika):

you're welcome :D

OpenStudy (anonymous):

yuppers